OpenStudy (anonymous):

Write an equation using slope point form that passes through the point (3,-6) and has a slope of 4.

OpenStudy (anonymous):

y=4x-6

OpenStudy (anonymous):

the work out is like this : y-y0=m(x-x0) y+6=4(x-3) y=4x-6

OpenStudy (mayaal):

y=-18x+4

OpenStudy (mayaal):

slope intercept equation:y=mx+b

OpenStudy (anonymous):